# LIFT: Improving Long Context Understanding of Large Language Models through Long Input Fine-Tuning

### v1

@@ 1 Introduction
Retrieval-free
Long-context adaptation-free
Large Language Models (LLMs), such as GPT-4 (Achiam et al., 2023), have revolutionized the field of natural language processing, driving breakthroughs in text generation and significant advancements in tasks like translation, summarization, and conversation. Lengthy sequences, which can span up to millions of tokens, are common in real-world applications including long books (Kočiskỳ et al., 2018), high-resolution videos (Wu et al., 2024; Tapaswi et al., 2016), and audio signals (Yang et al., 2024). Extending the context window allows models to capture dependencies across larger text spans and improve coherence, understanding, and accuracy in tasks that require reasoning over extended inputs.
However, as the context length increases, the computational complexity of the self-attention mechanism (Vaswani, 2017) grows quadratically, which limits the model’s ability to process long inputs. Additionally, storing a large number of attention weights and intermediate states like KV cache places a heavy burden on hardware resources. Moreover, it is challenging to capture long dependencies among pieces of information scattered throughout the long input while performing further comprehension and reasoning. Due to the limitation of context window, LLMs can hardly capture the overall information about a user’s query history or task input, resulting in suboptimal performance.
To address these challenges, researchers have developed various techniques to improve the long-context abilities of LLMs. Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Xu et al., 2023) and prompt compression (Jiang et al., 2023) aim to preprocess long inputs within the limited short context window by adaptive retrieval or text compression (El-Kassas et al., 2021). However, the effectiveness of these methods depends on the precision and relevance of the contextual information provided within the context window. It will lead to further hallucination when noisy, ambiguous, or conflicting information is provided. Long-context adaptation focuses on fine-tuning pretrained LLMs on corpora of long texts to extend their context windows (Chen et al., 2023b; Peng et al., 2023) and is more frequently used in more recent works. However, the adaptation process comes with significant costs in terms of both training data and computational resources. Additionally, with the extended context window, the cost of processing and generating long texts grows quadratically with the input length. Finally, despite the extension, the context window of these LLMs remains finite, preventing them from generalizing to inputs of infinite length.
@@
  • Efficient long-input training on the fly. LIFT dynamically adapts to newly introduced long inputs as fresh knowledge by adjusting model parameters, thereby eliminating the need for resource-intensive offline long-context adaptation. To enhance memorization of the long input, we segment it into overlapping pieces which can be fitted into a short context window and fine-tune the LLM on batches of the segments. Additionally, we improve long-context comprehension and reasoning through fine-tuning on well-designed auxiliary tasks, further optimizing performance on downstream applications.
  • Balancing in-parameter and in-context knowledge. As LIFT is mainly designed for short-context LLMs, the long input often need to be truncated to fit into their context windows. This necessitates a balance between leveraging the truncated in-context knowledge and the fine-tuned in-parameter knowledge. To address this issue, we propose a specialized attention adapter, Gated Memory, that automatically balances the long input memorization and comprehension in LIFT as well as the ICL ability of the original model.
  • Great improvement on popular long-context tasks. Our evaluations on several well-acknowledged long context benchmarks show that LIFT consistently benefits general tasks like long/short question answering (QA) and summarization across different base LLMs. For example, on the challenging long-dependency QA tasks of LooGLE (Li et al., 2023), the “LIFTed” LLama-3-8B model achieves an accuracy of 29.97%, significantly outperforming its pure ICL counterpart without LIFT which achieves only 15.44% accuracy.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG). RAG (Lewis et al., 2020) improves the performance of long-context understanding by integrating LLMs with external data sources or memory modules for retrieval (Xu et al., 2023; Jiang et al., 2024; Wang et al., 2024; Jin et al., 2024), thereby avoiding the need to process long inputs directly. Its performance heavily relies on the quality of the retrieved content, which must be relevant and concise enough to fit within the model’s short context window. RAG can experience significant performance degradation or hallucination issues when the retrieved context is inaccurate or mismatched.
Test-time training. Test-time training (TTT) (Liu et al., 2021; Gandelsman et al., 2022; Osowiechi et al., 2023; Hong et al., 2023) has emerged as a promising approach to adapt models to unseen data distributions during deployment, leveraging the test data to fine-tune the model at inference time. Recent works have applied similar ideas to improve model adaptability when dealing with lengthy, context-rich inputs (Sun et al., 2024; Behrouz et al., 2024), yet focus on proposing new architectures to replace Transformer and require pretraining from scratch. Our work, in contrast, introduces a continual learning perspective to the problem of long-context understanding, which focuses on improving arbitrary pretrained models’ long-context capabilities by fine-tuning the model on the long input, which is not restricted to specific models or layers.
@@ 3 Method
In this section, we introduce LIFT, a framework improving LLMs’ long context understanding through long input fine-tuning (Figure 1). The comparisons of our method with truncation, RAG, and the long context adaptation can be seen in Table 1 and implementation details are illustrated in Appendix A.
@@ 3.1 Training with Segmented Long Inputs
We propose a novel way to memorize the long inputs by storing them into LLMs’ parameters via fine-tuning. We formulize the memorization task as a language modeling task. Let the input be $\mathbf{x}=(x_{1},x_{2},\dots,x_{L})$, where $L$ is a very large number. The objective function for the task is defined as $\mathcal{L}_{LM}(\mathbf{x};\theta)=-\sum_{i=1}^{L}\log\mathbb{P}(x_{i}|\mathbf{x}_{1:i-1};\theta),$ where $\theta$ are the parameters.
Directly adapting the model to a piece of long text of length $L$ is challenging for LLMs whose context window lengths $l$ are shorter than $L$. Furthermore, it leads to a high computational complexity of $\mathcal{O}(L^{2})$ coping with such long inputs during training. One straightforward way is to cut $\mathbf{x}$ into $K$ non-overlapping short segments (trivial segmentation), denoted as $\mathbf{x}_{l_{1}:r_{1}},\dots,\mathbf{x}_{l_{K}:r_{K}}$, and fine-tune on batches of the segments. However, it fails to capture the sequentiality of the long input since the model cannot infer the correct order of the non-overlapping segments.
@@
where $l_{1}=1,r_{K}=L$, and $r_{k}-l_{k}+1=\ell,l_{k+1}=l_{k}+s,~{}\forall k=1,2,\dots,K-1$. Here $s$ is a hyperparameter controlling the overlap length of adjacent segments. Empirically, it is sufficient to use $s=\frac{3}{8}\ell$ leading to a constant computational overhead.
@@ 3.2 Training with Auxiliary Tasks
Adapting a pretrained LLM to a specific task poses the risk of impairing its other capabilities. Similarly, while adapting to the input helps the model memorize the input, it probably degrades other abilities, such as instruction-following. Moreover, effectively memorizing the long input doesn’t mean the model can reason based on it.
Adapting a pretrained LLM to a specific task poses the risk of impairing its other capabilities. Similarly, while adapting to the input may enhance the model’s ability to memorize it, this adaptation likely comes at the cost of other skills, such as instruction-following. Furthermore, successfully memorizing a lengthy input does not necessarily indicate that the model can reason effectively based on it.
To mitigate potential capability degradation while maintaining the reasoning capabilities of original model on the long context, we propose synthesizing auxiliary question-answering (QA) tasks, denoted as $(\mathbf{q}_{i},\mathbf{a}_{i})_{i=1}^{m}$, based on the long context. The objective function of the auxiliary tasks is defined as
@@ 3.3 Contextualized Training and Task Alignment
As discussed in Sections 3.1 and 3.2, we adapt an LLM to handle long inputs through two objectives: language modeling on segments of the long input and auxiliary QA tasks. While these tasks align with our objectives of memorizing the long input and enhancing reasoning based on the long input, the model may still struggle with the semantic divergence (memorization vs. reasoning) and structural divergence (language modeling vs. supervised fine-tuning) between different tasks. To address these challenges, we propose a contextualized training (CT) method for long input segments, shifting from the language modeling paradigm to a supervised fine-tuning paradigm and more closely aligning the task of input segment memorization and the auxiliary QA tasks.
@@
$\mathcal{L}_{input}(\mathbf{x};\theta)=-\sum_{k=1}^{K}\log\mathbb{P}(\mathbf{x}_{l_{k}:r_{k}}\!\mid\!concat(\mathbf{c},\mathbf{p});\theta),$ (4)
where $\mathbf{c}$ represents the given context, and $\mathbf{p}$ is a prompt instructing the model to recite the segment based on $\mathbf{c}$. For the QA tasks, we also modify the objective (2) by concatenating the questions with a context $\mathbf{c}_{q}$:
@@
where $\mathbf{c}_{q}$ keeps the same during training on different segments, which is only related to the test question. In this way, both the input memorization and QA tasks share a similar SFT format. In addition, they both align better with the real testing scenario, where given an LLM fine-tuned with LIFT, we can still fill the context window with the long input as much as possible to maximally leverage the in-context knowledge (or “working memory” of LLM), instead of only filling in the testing question. Such a technique greatly improves practical performance of LIFT.
To mitigate the risk of overfitting on the same $\mathbf{c}$, we further regularize $\mathbf{c}$ by randomly sampling $\mathbf{c}$ for each segment $\mathbf{x}_{l_{k}:r_{k}}$ from both the beginning and ending of the long input with a total length of $L$. Specifically, we select consecutive sentences from the beginning and ending respectively compositing $\mathbf{c}$ with a fixed length $l$ to align with the usages of contexts in real testing scenarios.
@@ 3.4 Gated Memory Architecture
The key intuition behind LIFT is to store parts of the long input that cannot fit into the context window directly in the model parameters. To achieve this, the adapter needs to effectively memorize these out-of-context parts and align its behavior with the scenario of having complete input. For a hypothetical complete input $(\mathbf{x}^{\prime},\mathbf{x})$ where $\mathbf{x}^{\prime}$ is the long input that we aim to absorb into model parameters and $\mathbf{x}$ represents the new in-context questions/prompts about the long input, we let their hidden states after the $(t-1)$-th layer be $({\hat{\mathbf{h}}}^{{}^{\prime}(t-1)},\hat{\mathbf{h}}^{(t-1)})$, where the length of $\mathbf{x}^{\prime}$ is $l^{\prime}$ and the length of $\mathbf{x}$ is $l$. We expect the following behavior in each layer of the original Transformer:
@@
$\operatorname{attn}(\hat{q}_{L},\hat{\mathbf{k}}_{1:L},\hat{\mathbf{q}}_{1:L})=\frac{\sum_{i=1}^{L}\exp(\langle\hat{q}_{L},\hat{k}_{i}\rangle)\hat{v}_{i}}{\sum_{j=1}^{L}\exp(\langle\hat{q}_{L},\hat{k}_{j}\rangle)},$ (6)
@@
where $\text{attn}(\hat{q}_{L},\hat{\mathbf{k}}_{l^{\prime}+1:L},\hat{\mathbf{v}}_{l^{\prime}+1:L})$ is the attention output with the same attention parameters operated on the in-context part’s hidden state $\hat{\mathbf{h}}^{(t-1)}$ (instead of the complete hidden state $({\hat{\mathbf{h}}}^{{}^{\prime}(t-1)},\hat{\mathbf{h}}^{(t-1)})$). Let $g$ and $m$ be implemented as neural networks. When the out-of-context input $\mathbf{x}^{\prime}$ is considered a constant and has been learned into the parameters of $g$ and $m$, $\hat{\mathbf{k}}_{1:l^{\prime}}$ and $\hat{\mathbf{v}}_{1:l^{\prime}}$ can be removed from $g$ and $m$. We further adopt an approximation to let $g$ only depend on $\hat{q}_{L}$. Consequently, both $g(\hat{q}_{L},\hat{\mathbf{k}}_{1:L})$ and $m(\hat{q}_{L},\hat{\mathbf{k}}_{1:L},\hat{\mathbf{v}}_{1:L})$ become functions of $\hat{q}_{L}$. The attention output simplifies to:
@@
The Gated Memory architecture is shown in Figure 3. In summary, we keep the original MLP and attention parameters (projectors for query, key, value and output) fixed and add two learnable MLPs for each attention head: 1) $g:\mathbb{R}^{d}\to\mathbb{R}$, a gating function that controls the ratio of information extracted from the memory and the context. 2) $m:\mathbb{R}^{d}\to\mathbb{R}^{d}$, a memory function which stores the out-of-context information of the long input and retrieves relevant memory based on the current query. In the ideal case where the learned $g$ and $m$ perfectly simulate Equations (7), that is, the previous information of $\hat{h}^{\prime}$ has been completely learned into the parameters of $g$ and $m$, the model can perform the same as the case with complete long input, thus achieving the effect of using a short-context model to simulate long-context transformers.

### v2

@@ 1 Introduction
Retrieval free
Long-context adaptation free
Large Language Models (LLMs), such as GPT-4 (Achiam et al., 2023), have revolutionized the field of natural language processing, driving breakthroughs in text generation and significant advancements in tasks like translation, summarization, and conversation. Long sequences, which can span up to millions of tokens, are common in real-world applications, including long books (Kočiskỳ et al., 2018), accounting documentsdocumentsdocuments (Li et al., 2024), high-resolution videos (Wu et al., 2024; Tapaswi et al., 2016), and audio signals (Yang et al., 2024). Extending context windows allows models to capture dependencies across larger text spans and improve coherence, understanding, and accuracy in tasks that require reasoning over extended inputs.
However, as context lengths increase, the computational complexity of the self-attention mechanism (Vaswani, 2017) grows quadratically, which limits models’ ability to process long inputs. Additionally, storing a large number of attention weights and intermediate states like KV cache places a heavy burden on hardware resources. Moreover, it is challenging to capture long dependencies among pieces of information scattered throughout long inputs while performing further comprehension and reasoning. Due to the limitation of context windows, LLMs can hardly capture the overall information about a user’s query history or task input, resulting in suboptimal performance.
To address these challenges, researchers have developed various techniques to improve the long-context abilities of LLMs. Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Xu et al., 2023) and prompt compression (Jiang et al., 2023) aim to preprocess long inputs within a limited short context window by adaptive retrieval or text compression (El-Kassas et al., 2021). However, the effectiveness of these methods depends on the precision and relevance of the contextual information provided within the context window. It will lead to further hallucinations when noisy, ambiguous, or conflicting information is provided. Long-context adaptation focuses on fine-tuning pretrained LLMs on corpora of long texts to extend their context windows (Chen et al., 2023b; Peng et al., 2023) and is more frequently used in more recent works. However, the adaptation process comes with significant costs in terms of both training data and computational resources. Additionally, with the extended context window, the cost of processing and generating long texts grows quadratically with the input length. Finally, despite the extension, the context windows of these LLMs remain finite, preventing them from generalizing to inputs of infinite length.
@@
  • Efficient long-input training on the fly. LIFT dynamically adapts to newly introduced long inputs as fresh knowledge by adjusting model parameters, thereby eliminating the need for resource-intensive offline long-context adaptation. To enhance memorization of the long input, we segment it into overlapping segments which can be fitted into a short context window and fine-tune the LLM on batches of the segments. Additionally, we improve long-context comprehension and reasoning through fine-tuning on well-designed auxiliary tasks, further optimizing performance on downstream applications.
  • Balancing in-parameter and in-context knowledge. As LIFT is mainly designed for short-context LLMs, the long input often needs to be truncated to fit into their context windows. This necessitates a balance between leveraging the truncated in-context knowledge and the fine-tuned in-parameter knowledge. To address this issue, we propose a specialized attention adapter, Gated Memory, that automatically balances the long input memorization and comprehension in LIFT as well as the ICL ability of the original model.
  • Great improvement on popular long-context tasks. Our evaluations on several well-acknowledged long context benchmarks show that LIFT consistently benefits general tasks like long/short question answering (QA) and summarization across different base LLMs. For example, on the challenging long-dependency QA tasks of LooGLE (Li et al., 2023), the “LIFTed” Llama-3-8B-Instruct model achieves an accuracy of 29.97%, significantly outperforming its pure ICL counterpart without LIFT which achieves only 15.44% accuracy.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG). RAG (Lewis et al., 2020) improves the performance of long-context understanding by integrating LLMs with external data sources or memory modules for retrieval (Xu et al., 2023; Jiang et al., 2024; Wang et al., 2024; Jin et al., 2024), thereby avoiding the need to process long inputs directly. Its performance heavily relies on the quality of retrieved content, which must be relevant and concise enough to fit within models’ short context windows. RAG can experience significant performance degradation or hallucination issues when the retrieved context is inaccurate or mismatched.
Test-time training. Test-time training (TTT) (Liu et al., 2021; Gandelsman et al., 2022; Osowiechi et al., 2023; Hong et al., 2023) has emerged as a promising approach to adapt models to unseen data distributions during deployment, leveraging test data to fine-tune the model at inference time. Recent works have applied similar ideas to improve model adaptability when dealing with lengthy, context-rich inputs (Sun et al., 2024; Behrouz et al., 2024), yet focus on proposing new architectures to replace Transformer and require pretraining from scratch. Our work, in contrast, introduces a continual learning perspective to the problem of long-context understanding, which focuses on improving arbitrary pretrained models’ long-context capabilities by fine-tuning them on the long input, which is not restricted to specific models or layers.
@@ 3 Method
In this section, we introduce LIFT, a framework improving LLMs’ long context understanding through long input fine-tuning (Figure 1). The comparisons of our method with RAG and long-context adaptation can be seen in Table 1 and the implementation details are illustrated in Appendix A.
@@ 3.1 Training with Segmented Long Inputs
We propose a novel way to memorize long inputs by storing them into LLMs’ parameters via fine-tuning. We formulize the memorization task as a language modeling task. Let the input be $\mathbf{x}=(x_{1},x_{2},\dots,x_{L})$, where the input length $L$ is a very large number. The objective function for the task is defined as $\mathcal{L}_{LM}(\mathbf{x};\theta)=-\sum_{i=1}^{L}\log\mathbb{P}(x_{i}|\mathbf{x}_{1:i-1};\theta),$ where $\theta$ are the parameters.
Directly adapting a model to a piece of long text of length $L$ is challenging for LLMs whose context window lengths $l$ are shorter than $L$. Furthermore, it leads to a high computational complexity of $\mathcal{O}(L^{2})$ coping with such long inputs during training. One straightforward way is to cut $\mathbf{x}$ into $K$ non-overlapping short segments (trivial segmentation), denoted as $\mathbf{x}_{l_{1}:r_{1}},\dots,\mathbf{x}_{l_{K}:r_{K}}$, and fine-tune the model on batches of the segments. However, trivial segmentation fails to capture the sequentiality of the long input since the model cannot infer the correct order of the non-overlapping segments.
@@
where $l_{1}=1,r_{K}=L$, and $r_{k}-l_{k}+1=\ell,l_{k+1}=l_{k}+s,~{}\forall k=1,2,\dots,K-1$. Here $s$ is a hyperparameter controlling the overlap length of adjacent segments. Empirically, it is sufficient to use $s=\frac{3}{8}\ell$, leading to a constant computational overhead.
@@ 3.2 Training with Auxiliary Tasks
Adapting a pretrained LLM to a specific task poses the risk of impairing its other capabilities. Similarly, while adapting to the input helps the model memorize the input, it probably degrades other abilities, such as instruction-following. Moreover, successfully memorizing the lengthy input does not necessarily indicate that the model can reason effectively based on it.
To mitigate potential capability degradation while maintaining the reasoning capabilities of original model on the long context, we propose synthesizing auxiliary question-answering (QA) tasks, denoted as $(\mathbf{q}_{i},\mathbf{a}_{i})_{i=1}^{m}$, where $m$ is the number of auxiliary tasks, based on the long input. The objective function of the auxiliary tasks is defined as
@@ 3.3 Contextualized Training and Task Alignment
As discussed in Sections 3.1 and 3.2, we adapt an LLM to handle a long input through two objectives: language modeling on segments of the long input and auxiliary QA tasks. While these tasks align with our objectives of memorizing the long input and enhancing reasoning based on the long input, the model may still struggle with the semantic divergence (memorization vs. reasoning) and structural divergence (language modeling vs. supervised fine-tuning) between different tasks. To address these challenges, we propose a contextualized training (CT) method for long input segments, shifting from the language modeling paradigm to a supervised fine-tuning paradigm and more closely aligning the task of input segment memorization and the auxiliary QA tasks.
@@
$\mathcal{L}_{input}(\mathbf{x};\theta)=-\sum_{k=1}^{K}\log\mathbb{P}(\mathbf{x}_{l_{k}:r_{k}}\!\mid\!concat(\mathbf{c}_{k},\mathbf{p});\theta),$ (4)
where $\mathbf{c}_{k}$ represents the given context, and $\mathbf{p}$ is a prompt instructing the model to recite the segment based on $\mathbf{c}_{k}$. For the QA tasks, we also modify the objective (2) by concatenating the questions with a context $\mathbf{c}_{q}$:
@@
where $\mathbf{c}_{q}$ keeps the same during training on different segments, which is only related to the test question. In this way, both the input memorization and QA tasks share a similar SFT format. In addition, they both align better with the real testing scenario, where given a LIFTed LLM, we can still fill the context window with the long input as much as possible to maximally leverage the in-context knowledge, instead of only filling in the testing question. Such a technique greatly improves practical performance of LIFT.
To mitigate the risk of overfitting, instead of using the same $\mathbf{c}_{k}$ for all the segments $\mathbf{x}_{l_{k}:r_{k}}$, we further regularize $\mathbf{c}_{k}$ by randomly sampling $\mathbf{c}_{k}$ for each segment $\mathbf{x}_{l_{k}:r_{k}}$ from both the beginning and ending of the long input with a total length of $L$. Specifically, we select consecutive sentences from the beginning and ending respectively compositing $\mathbf{c}_{k}$ with a fixed length $l$ to align with the usages of contexts in real testing scenarios.
@@ 3.4 Gated Memory Architecture
The key intuition behind LIFT is to store the parts of a long input that cannot fit into the context window directly in model parameters. To achieve this, the adapter needs to effectively memorize these out-of-context parts and align its behavior with the scenario of having complete input. For a hypothetical complete input $(\mathbf{x}^{\prime},\mathbf{x})$ where $\mathbf{x}^{\prime}$ is the long input that we aim to absorb into model parameters and $\mathbf{x}$ represents the new in-context questions/prompts about the long input, we let their hidden states after the $(t-1)$-th layer be $({\hat{\mathbf{h}}}^{{}^{\prime}(t-1)},\hat{\mathbf{h}}^{(t-1)})$, where the length of $\mathbf{x}^{\prime}$ is $l^{\prime}$ and the length of $\mathbf{x}$ is $l$. In practice, the model has access to the questions/prompts $\mathbf{x}$ only and we let their hidden states after the $(t-1)$-th layer be $\mathbf{h}^{(t-1)}$. We expect the following behavior in each layer of the original Transformer:
@@
$\operatorname{attn}(\hat{q}_{L},\hat{\mathbf{k}}_{1:L},\hat{\mathbf{v}}_{1:L})=\frac{\sum_{i=1}^{L}\exp(\langle\hat{q}_{L},\hat{k}_{i}\rangle)\hat{v}_{i}}{\sum_{j=1}^{L}\exp(\langle\hat{q}_{L},\hat{k}_{j}\rangle)},$ (6)
@@
where $\text{attn}(\hat{q}_{L},\hat{\mathbf{k}}_{l^{\prime}+1:L},\hat{\mathbf{v}}_{l^{\prime}+1:L})$ is the attention output with the same attention parameters operated on the in-context part’s hidden state $\hat{\mathbf{h}}^{(t-1)}$ (instead of the complete hidden state $({\hat{\mathbf{h}}}^{{}^{\prime}(t-1)},\hat{\mathbf{h}}^{(t-1)})$). Let $g$ and $m$ be implemented as neural networks. When the out-of-context input $\mathbf{x}^{\prime}$ is considered a constant and has been absorbed into the parameters of $g$ and $m$, $\hat{\mathbf{k}}_{1:l^{\prime}}$ and $\hat{\mathbf{v}}_{1:l^{\prime}}$ can be removed from $g$ and $m$. We further adopt an approximation to let $g$ only depend on $\hat{q}_{L}$. Consequently, both $g(\hat{q}_{L},\hat{\mathbf{k}}_{1:L})$ and $m(\hat{q}_{L},\hat{\mathbf{k}}_{1:L},\hat{\mathbf{v}}_{1:L})$ become functions of $\hat{q}_{L}$. The attention output simplifies to:
@@
The Gated Memory architecture is shown in Figure 3. In summary, we keep the original MLP and attention parameters (projectors for query, key, value and output) fixed and add two learnable MLPs for each attention head: 1) $g:\mathbb{R}^{d}\to\mathbb{R}$, a gating function that controls the ratio of information extracted from the memory and the context. 2) $m:\mathbb{R}^{d}\to\mathbb{R}^{d}$, a memory function which stores the out-of-context information of the long input and retrieves relevant memory based on the current query. In the ideal case where the learned $g$ and $m$ perfectly simulate Equations (7), that is, the previous information of $\hat{h}^{\prime}$ has been completely absorbed into the parameters of $g$ and $m$, the model can perform the same as the case with complete long input, thus achieving the effect of using a short-context model to simulate long-context transformers.